OpenStudy (anonymous):

Please help me with my Algebra II Consider the function f (x) = (1/3)x + 2 a. Find the inverse of f(x) and name it g(x). Show and explain your work. b. Use function composition to show that f(x) and g(x) are inverses of each other.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

The inverse of a function is found by interchanging x with y, and y with x, AND then solving for y. Since f(x) = (1/3)x + 2, that means y = (1/3)x + 2 The inverse is x = (1/3)y + 2 Now, solving for y, we have x - 2 = (1/3)y Now multiply both sides by 3, we get y = 3(x-2) or 3x - 6 So the inverse is 3x - 6. We say g(x) = 3x - 6

OpenStudy (anonymous):

So was that all for A? I am soo confused!

OpenStudy (anonymous):

yes. What are you confused about?

OpenStudy (anonymous):

I think I got it after looking through that for a while... Now, How does function composition work? My teacher wasn't the greatest at describing this lesson.

OpenStudy (anonymous):

I managed to figure out the rest! Thanks!